OpenStudy (anonymous):

Medal and will become a fan. What is the general form of the equation of a circle with a center at (–5, 0) and a radius of 4?

OpenStudy (anonymous):

do you know general form of a circle?

OpenStudy (anonymous):

x^2 + y^2 + Ax + By + C = 0

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

the center is ab.

OpenStudy (anonymous):

so the center is A=-5 and B=0 right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

so lets put it in standard form first. (x-a)^2 + (y-b)^2 = r^2

OpenStudy (anonymous):

(x+5)^2 + (y-0)^2= 4^2

OpenStudy (anonymous):

now expand.

OpenStudy (anonymous):